OpenStudy (anonymous):

If two rulers are chosen randomly one at a time from the box, without replacing the first one, what is the theoretical probability that both rulers are yellow? 1/25 Red 5 Blue 2 Green 4 Yellow 4 Orange 5 *1/5 x 1/5 =1/25 ?

Parth (parthkohli):

Let us start by counting the total number of rulers - 5 + 2 + 4 + 4 + 5 = 20 P(1st time yellow) = 4/20 = 1/5 Now that we won't replace, then it's one less from both the numerator AND denominator. P(2nd time yellow) = 3/19 \(\Large \color{MidnightBlue}{\Rightarrow {1 \over 5} * {3 \over 19} = {3 \over 95} }\)

OpenStudy (anonymous):

well no shame in trying thanks

Parth (parthkohli):

Absolutely lol